Assume a population of 1,000 people of when 15 are sick with a certain disease, 12 of the 15 die from this disease.

Calculate the mortality rate from the disease.

Supposed you're talking to a 5 year old. Explain difference between prevalence and incidence.

Active surveillance is usually more accurate than passive surveillance.  

True

False

In 2006 there were 5,600 deaths due to cancer in a certain community of individuals aged 40-65 years. The community has a population of 45,000. The expected number of deaths in the community, based on age-specific death rates for cancer was 6,700 during 2006. What was the standardized mortality ratio (SMR) for cancer in this community?

.84

8.04

1.2

.12

Can’t be determined
Answer #1

Mortality rate = total number of population/ actual death.  

1000/12=83.3

Prevalence: it is proportion of particular population found to be affected by a medical condition.

Incidence: accurate rate of some thing happens.

True..Active surveillance is usually more accurate than passive surveilance..

Standardized mortality ratio we can get by total number of population divided with actual number of death

SMR= 45,000/5600 = 8.04